QUESTION: SOLVE THE SYSTEM BY SUBSTIUTION
2X+2Y=0
Y=4X+0

Answer by Lacey020991(49)

Where the y is on the first equation, your going to the second equation in and solve.
.
2x + 2(4x+0) = 0 ~~~~~Distribute the 2
2x + 8x + 0 = 0 ~~~~~~Combine like terms
10x = 0 ~~~~~~~~~~~~~~Divide both sides by 10
x = 0
.
Then plug the x value back into one of the equations:
y = 4(0) + 0
y = 0
.
So your answer would be (0,0)
They intersect at the origin